G21 imate afirmstree cash flows and returns Fill in the blue color cells S Usethe given inputs in the previous worksheet to do thecalculations 6 Turn in a hard copy of this worksheet when you cometo the exam. No need to turn in the inputs worksheet. 9 Firms perspective 2017 8,868 2,217) EBIT(1-t) change in fixed assets change in noncash WC 12 13 Cash flow to firm FCFF FCFF (as % of EBIT(1-t)) 68% 15 16 Return 17 ROC 7.48%) use average capital 19 Equitys perspective 2017 Net Income 21 change in fxed assets change in noncash WC +increase in debt FCFE FCFE (as % of Net Income) 23 24 Cash flow to equity 25 26 27 Return 28 29 ROE use average eqwtyCompany name Company ticker in unit of Comcast Corporation CMCSA million Year 2017 2016 Operating Income or LoSS, or EBIT tax rate (- tax/income before tax) 17,545 49.46% 16,920 9 77% Net Income 22,714 8,695 Receivables Inventory Accounts Payable noncash WC 8,546 7,955 6,926 1,620 6,915 1,040 Fixed assets, or net PPE 38,470 36,253 Short-term Debt Long-term Debt Total Debt 64,556 59,100 123,656 61,046 55,566 116,612 Equity 68,606 53,943 Capital 127,706 109,509
Answer #1
A B C D E F G H I J
2
3 Free cash flow to the firm (FCFF) can be calculated as follows:
4 Free Cash Flow = EBIT*(1-Tax Rate)+Depreciation - Capital Expenditures - Change in working capital
5
6 ROC =NOPAT / Average capital
7 NOPAT =EBIT*(1-Tax Rate)
8
9 FCFE =Net Income - Capital Expenditures - Change in working capital + Increase in Debt
10 ROE =Net Income / Average Equity
11
12 Firm's Perspective Year 2017
13 EBIT(1-t) $8,867.24 =17545*(1-49.46%)
14 -Change in fixed asset ($2,217.00) =-(38470-36253)
15 -change in noncash WC ($580.00) =-(1620-1040)
16 Cash flow to firm FCFF $6,070.24 =SUM(E13:E15)
17 FCFF (as % of EBIT*(1-t)) 68.46% =E16/E13
18
19 Return ROC 7.48% =E13/AVERAGE(127706,109509)
20
21 Equity Perspective Year 2017
22 Net Income $22,714.00
23 -Change in fixed asset ($2,217.00) =-(38470-36253)
24 -change in noncash WC ($580.00) =-(1620-1040)
25 + increase in debt $7,044.00 =(123656-116612)
26 Cash flow to equity FCFE $19,917.00 =SUM(E22:E24)
27 FCFE (as % of Net Income) 87.69% =E26/E22
28
29 Return ROE 37.07% =E22/AVERAGE(68606,53943)
30